OpenStudy (anonymous):

Indicate in standard form the equation of the line passing through the given point and having the given slope. D(5, -2) m = 2/5

OpenStudy (anonymous):

Eq. of a line thro' (x1,y1) with slope m is y-y1=m(x-x1) y-(-2)=2/5 (x-5) 5(y+2)=2(x-5) 5y+10=2x-10 or 5y=2x-10-10 or 5y=2x-20

OpenStudy (anonymous):

perfect thats what i got thanks